OpenStudy (anonymous):

If there are 25,000 units of heparin in 500ml, how many units of heparin are there in 176ml?

OpenStudy (saifoo.khan):

easy medal.

OpenStudy (saifoo.khan):

8750

OpenStudy (saifoo.khan):

25000 = 500 x = 176 cross multiply.

OpenStudy (anonymous):

25,000 units of heparin in 500ml units of heparin in 1ml=25000/500 units of heparin in 176ml=(25000/500)*176=8800units

OpenStudy (anonymous):

Awesome!!! Thanks